OpenStudy (flowers):

Can someone help me with composing two polynomial functions? F(x)=x^2+4. g(x)=x-5. (f*g)(x)=?

OpenStudy (danjs):

the product of two functions. (f * g)(x) = f(x) * g(x)

OpenStudy (sunnnystrong):

What @DanJS said... f composed with g --> \[f(x-5)\] \[(x-5)^2+4\]

OpenStudy (flowers):

Oh okay basically substitutions I dont like functions makes my brain crazy

OpenStudy (danjs):

yeah , just multiply them together

OpenStudy (sunnnystrong):

... wait, is this supposed to be f circle g or f*g?

OpenStudy (danjs):

(f*g)(x) = f(x) * g(x) = (x^2 + 4)*(x - 5) = ...

OpenStudy (flowers):

Circle

OpenStudy (danjs):

thats not multiply, open circle is composition function f(g(x))

OpenStudy (sunnnystrong):

Oh okay cool. Yep you just compose your function --> If it was g circle f what would it be?

OpenStudy (sunnnystrong):

yea idk @DanJS XD

OpenStudy (flowers):

Oops i didnt notice a circle button my computer

OpenStudy (sunnnystrong):

DW about it hahaha... but what would g composed with f look like? @flowers

OpenStudy (danjs):

the f is first, it is 'f of g' f[g(x)]

OpenStudy (flowers):

Im not sure but X^2+4-x-5

OpenStudy (sunnnystrong):

@DanJS is right... so this is different. (f circle g does not equal g circle f ) just to be clear... It is all basically substitution like you said :D ^^ See earlier comment @flowers If: F(x)=x^2+4. G(x)=x-5. G(x^(2)+4) (x^(2)+4)-5 =x^(2)-1

OpenStudy (flowers):

Cool thanks Thanks for your help

OpenStudy (sunnnystrong):

NP :D Always happy to help!